Calculate the amount of simple interest earned:Principal: $2500Annual Rate: 2%TIme: 4.5 years
Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

Amount of Interest: P*r*t = 2500*0.02*1 = Use your calculator.

Answer by MathTherapy(10552)


Calculate the amount of simple interest earned:Principal: $2500Annual Rate: 2%TIme: 4.5 years
Simple interest formula: PRT, so MULTIPLY them.
Hint: It's NOT 2,500 * 1 * .02